Find the prime factorization of three common multiples of 7560 and 18000.

Answer by stanbon(75887)

Find the prime factorization of three common multiples of 7560 and 18000.
----
7560/2 = 3780
3780/2 = 1890
1890/2 = 945
945/3 = 315
945/3 = 105
105/3 = 35
35/5 = 7
----
Prime factor form: 5*7*3^3*2^3
====================================
Use the same process to find the answer for 18000.
Cheers,
Stan H.
================